OpenStudy (anonymous):

x^3+125; (x+5) find the other factor

OpenStudy (whpalmer4):

We can write any polynomial as a product of factors. That means \[x^3+125 = (x+5)(\text{other factors})\]Do you know how to divide polynomials, either by long division or synthetic division?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

isn't there another way to solve it without division?

OpenStudy (whpalmer4):

Well, it works pretty much the same way it does for numbers. What's the biggest power of \(x\) we can multiply with \((x+5)\) and still be able to subtract it from \(x^3+125\)?

OpenStudy (anonymous):

2?

OpenStudy (whpalmer4):

We've got \(x\), and we want to turn it into \(x^3\), so we multiply by \(x^2\): \[x^2(x+5) = x^3 + 5x^2\] Now we subtract that from our original polynomial: \[x^3+125 - (x^3+5x^2) = -5x^2 + 125\]So our remainder after 1 division step is \[-5x^2+125\] What's the biggest thing we can multiply with \((x+5)\) for the next division step? We want to get a \(-5x^2\) term...

OpenStudy (anonymous):

the x^3 cancels out right?

OpenStudy (whpalmer4):

We've already taken out the x^3 term, yes. now we are trying to divide \(-5x^2+125\) by \(x+5\)

OpenStudy (anonymous):

how did u get -5x^2?

OpenStudy (whpalmer4):

Here, maybe drawing it in the usual fashion will help: x^2 ----------------------------- x+5 | x^3 + 0x^2 + 0x + 125 x^3 + 5x^2 -------------------- 0x^3 - 5x^2 + 0x + 125

OpenStudy (anonymous):

are you doing synthetic substitution?

OpenStudy (whpalmer4):

now we're looking to subtract off the \(-5x^2\) term, so we multiply \((-5x)(x+5) = -5x^2-25x\) and subtract that as our next step x^2 -5x ----------------------------- x+5 | x^3 + 0x^2 + 0x + 125 x^3 + 5x^2 -------------------- 0x^3 - 5x^2 + 0x + 125 -5x^2 -25x ---------------- 0x^2 + 25x + 125 Now, what can we multiply with \(x+5\) to get \(25x + 125\)?

OpenStudy (whpalmer4):

no, this is long division

OpenStudy (whpalmer4):

so what do you have to multiply with (x+5) to get 25x + 125?

OpenStudy (anonymous):

5x+25

OpenStudy (anonymous):

right?

OpenStudy (whpalmer4):

no, not quite. (5x+25)(x+5) = 5x^2 + 25x + 25x + 125 how about just 25? 25(x+5) = 25x + 5*25 = 25x + 125

OpenStudy (whpalmer4):

so that means \[\frac{x^3+125}{x+5} = x^2-5x+25\] and we can factor \[x^3+125 = (x+5)(x^2-5x+25)\] Can we factor \(x^2-5x+25\) some more?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i put that into the calculator and i just got 25, @whpalmer4

OpenStudy (whpalmer4):

you put what into the calculator?

OpenStudy (anonymous):

no i get it now, its okay. thanks

OpenStudy (whpalmer4):

we can't factor\( x^2-5x+25\) any more because the only factors we have to work with are either 1 and 25, or 5 and 5, and there's no combination of them that we could get to multiply to 25 and sum to -5. Therefore, it is fully factored.